I am doing this problem and I am getting the z value correctly, but I cannot figure out the p value. I am using a calculator so I anyone can explain how to find the p value for me that would be great.

Assume that z is the test statistic.

(a) Ho: μ = 22.5, Ha: μ > 22.5; x = 26.4, σ = 7.1, n = 20
(i) Calculate the test statistic z. (Give your answer correct to two decimal places.)
z= 2.46

(ii) Calculate the p-value. (Give your answer correct to four decimal places.)


(b) Ho: μ = 200, Ha: μ < 200; x = 192.9, σ = 35, n = 25
(i) Calculate the test statistic z. (Give your answer correct to two decimal places.)

z= -1.01

(ii) Calculate the p-value. (Give your answer correct to four decimal places.)
Answer #1

Solution:

a) ii)

This is a right tailed test,

P-value = P(Z > z)

= 1 - P(Z < 2.46 )

= 1 - 0.9931

= 0.0069

b) ii)

This is a left tailed test.

P-value = P(Z < z )

= P(Z < -1.01 )

= 0.1562